OpenStudy (anonymous):

-10√18 - 5√32 simplify

OpenStudy (anonymous):

-1550

OpenStudy (anonymous):

how do u work it out

OpenStudy (anonymous):

-1018 - 532

OpenStudy (angela210793):

-10*3sqr2-5*4sqr2=-30sqr2-20sqr2=-50sqr2

OpenStudy (anonymous):

woah...i read the question wrong again

OpenStudy (anonymous):

\[-10\sqrt{9*2}-5\sqrt{2*16}\]

OpenStudy (angela210793):

express sqr18 as sqr(9*2)=3sqr2 express sqr32as sqr(16*2)=4sqr2

OpenStudy (angela210793):

then i guess y understand

OpenStudy (anonymous):

\[-30\sqrt{2}-20\sqrt{2}\] \[-50\sqrt{2}\]